### 06 Mar 2024 v2

# A Novel Image Classification Framework Based on Variational Quantum Algorithms

Abstract
Image classification is a crucial task in machine learning with widespread practical applications. The existing classical framework for image classification typically utilizes a global pooling operation at the end of the network to reduce computational complexity and mitigate overfitting. However, this operation often results in a significant loss of information, which can affect the performance of classification models. To overcome this limitation, we introduce a novel image classification framework that leverages variational quantum algorithms (VQAs)—hybrid approaches combining quantum and classical computing paradigms within quantum machine learning. The major advantage of our framework is the elimination of the need for the global pooling operation at the end of the network. In this way, our approach preserves more discriminative features and fine-grained details in the images, which enhances classification performance. Additionally, employing VQAs enables our framework to have fewer parameters than the classical framework, even in the absence of global pooling, which makes it more advantageous in preventing overfitting. We apply our method to different state-of-the-art image classification models and demonstrate the superiority of the proposed quantum architecture over its classical counterpart through a series of experiments on public datasets. Our experiments show that the proposed quantum framework achieves up to a 9.21% increase in accuracy and up to a 15.79% improvement in F1 score, compared to the classical framework.
1 Introduction
In light of recent breakthroughs in quantum technologies, particularly the availability of noisy intermediate-scale quantum (NISQ) processors [cheng2023noisy], the field of quantum machine learning has attracted growing concerns and triggered an enormous amount of work [schuld2015introduction, biamonte2017quantum, dunjko2018machine, zhang2020recent, cerezo2022challenges]. Quantum machine learning is an interdisciplinary field that combines concepts from quantum physics and machine learning to develop innovative algorithms and models. By harnessing the power of quantum computing, quantum machine learning aims to improve the performance of machine learning algorithms. While still an emerging discipline, quantum machine learning has already demonstrated promising quantum extensions to classical machine learning techniques including support vector machines [varatharajan2018big], clustering [kerenidis2018q, otterbach2017unsupervised], and principal component analysis [article].
Image classification is a fundamental task in computer vision that involves categorizing images into predefined classes or categories. Over the past few years, this field has witnessed rapid progress. Convolutional neural networks (CNNs) emerged as a breakthrough technique, significantly improving image classification accuracy. Various CNN architectures [krizhevsky2012imagenet, lin2013network, simonyan2014very, szegedy2015going, he2016deep, huang2017densely, tan2019efficientnet] have progressively advanced state-of-the-art (SOTA) results on benchmark datasets such as ImageNet [deng2009imagenet]. Meanwhile, self-attention models like Transformers [vaswani2017attention] in natural language processing have been introduced to computer vision. The corresponding transformer-based vision models, such as Vision Transformer (ViT) [dosovitskiy2010image] and Swin Transformer [liu2021swin], can achieve compelling results on image recognition and even outperform CNNs on some tasks. More recently, hybrid approaches combining CNN and transformer modules are also gaining attention. These hybrid CNN-transformer models, including CoAtNet [dai2021coatnet] and MaxViT [tu2022maxvit], have been demonstrated to improve the generalizability of pure transformer-based models and achieve excellent performance in image classification without being pre-trained on large-scale image datasets.
In most image classification architectures, the input images are processed by a backbone model, also referred to as a feature extractor, which extracts relevant features from the input. This backbone model generally consists of a series of convolutional layers, transformer-based layers, or a combination of both. The resulting feature maps are typically fed into a global pooling layer followed by fully connected layers and classifier. Global pooling is a commonly employed technique in the field of image recognition [lin2013network, simonyan2014very, szegedy2015going, he2016deep, huang2017densely, tan2019efficientnet, liu2021swin, dai2021coatnet, tu2022maxvit]. It aggregates the spatial information of feature maps across the entire image into a single vector representation. The advantage of global pooling lies in its ability to capture the overall context and semantic information of an image, resulting in a compact and fixed-length representation that can be easily fed into subsequent layers for classification tasks. Global pooling also helps to mitigate the spatial variance of features, enabling the model to achieve translation invariance and robustness to object deformations. Moreover, it reduces the dimensionality of the feature maps, leading to computational efficiency and alleviating the risk of overfitting.
While global pooling provides some benefits for image classification, it has several limitations. Collapsing an entire feature map into a single vector results in the complete loss of fine-grained spatial information. The compact fixed-length representation has also restricted expressive capacity compared to a fully-connected layer operating on spatial data, which may hinder capturing complex spatial relationships. This can be disadvantageous for tasks that require precise object localization, discriminative information or detailed spatial information. Furthermore, uninformative features corresponding to background or irrelevant regions can dominate the global summary statistic and render it unresponsive to salient features. Additionally, since global pooling treats the entire image equally, it may not effectively capture local variations or small-scale patterns that are important for accurate recognition.
Researchers have explored different strategies to enhance the performance of global pooling and mitigate its drawbacks. Firstly, several variants of global pooling are proposed to provide a trade-off between the global average pooling and global max pooling with a nonlinear log-average-exp (LAE) [zhang2018alphamex, lowe2021logavgexp] or log-sum-exp (LSE) [sun2016pronet] function, which helps extract features more effectively. Another approach is to leverage pyramid pooling [he2015spatial, jose2018pyramid, qi2018concentric, lin2019fourier] as an alternative to global pooling. Pyramid pooling methods divide the image into multiple regions of different scales and extracts features from each region independently. This allows for the capture of multi-scale information and the preservation of spatial details. Moreover, attention mechanisms have also been integrated with global pooling to improve its discriminative power [qiu2018global, zhang2020global, behera2021context]. These mechanisms selectively emphasize salient regions or features while suppressing irrelevant ones, enhancing the model’s ability to capture important information. In addition, spatial transformer networks (STNs) [jaderberg2015spatial] have been proposed to enable adaptive spatial transformations prior to pooling. STNs apply learned transformations to align features, enhancing the localization accuracy of global pooling. Furthermore, the method of global second-order pooling [li2018towards, gao2019global, wang2020deep, song2022eigenvalues] is also exploited for capturing more discriminative image representations. This approach calculates covariance matrices to obtain higher-order statistical information between features.
In contrast to the above methods, our work employs quantum machine learning algorithms to address the issue of global pooling for the task of image classification. Quantum machine learning has the potential to revolutionize the area of image classification as quantum algorithms can process large datasets of images more efficiently than classical algorithms, leading to faster and more accurate classification [dang2018image, kerenidis2019quantum, mari2020transfer, henderson2020quanvolutional, li2020quantum, liu2021hybrid, henderson2021methods, houssein2021hybrid, sagingalieva2022hyperparameter, hur2022quantum, senokosov2023quantum, melnikov2023quantum]. In this paper, we propose a novel image classification framework based on variational quantum algorithms (VQAs) [mcclean2016theory, bharti2022noisy, cerezo2021variational] in the context of quantum machine learning. VQAs are a class of quantum algorithms that utilize classical computers to optimize a parameterized quantum circuit. Due to their hybrid quantum-classical approach that has potential noise resilience [mcclean2016theory], VQAs have been considered as one of the leading candidates to achieve application-oriented quantum computational advantage on NISQ devices.
The proposed framework differs from classical architectures primarily in that global pooling operations at the end of the network are no longer required. This is due to the introduction of the variational quantum circuits (VQCs) [cerezo2021variational, mitarai2018quantum, farhi2018classification] which possess remarkable data storage capacity. Consequently, our framework can fully utilize the feature maps extracted via the backbone model for image classification, which is particularly advantageous for tasks requiring fine-grained features. Moreover, despite the removal of the global pooling layer, our framework does not suffer from the issue of excessive model parameters and overfitting, as commonly encountered in classical deep learning frameworks. In contrast, our framework actually has fewer parameters, thanks to the robust data storage and expressive capabilities of the variational quantum circuits. Furthermore, our framework exhibits high flexibility and adaptability, making it applicable to most of existing deep learning models for image classification.
In summary, the contributions of our work are
  • We propose a novel image classification framework based on variational quantum algorithms, which does not require global pooling before the classification layer and enables models to capture more discriminative details from images using fewer parameters. To the best of our knowledge, our work is the first attempt to tackle the problem of global pooling in the classical image classification framework by employing quantum machine learning algorithms.
  • We evaluate our proposed framework on four challenging datasets, namely Croatian Fish [jager2015croatian], Aircraft [maji2013fine], Breast Ultrasound image [al2020dataset], and Apples or Tomatoes [AOT]. Experimental results demonstrate that our framework significantly outperform the classical framework.
2 Method
2.1 Preliminaries
2.1.1 Global Pooling
Pooling is a common operation in image classification models that plays a vital role in reducing spatial dimensions and capturing invariant features. It aims to down sampling feature maps by summarizing the presence of features in patches of the feature map. Two commonly used pooling methods are average pooling and max pooling. Average pooling calculates the average value of the feature within each pooling region while max pooling selects the maximum value of the feature. Examples of these two pooling approaches are shown in Fig. 1.
Figure 1: Examples of two types of pooling operations. (a) Average pooling with a pooling area of size 2x2 and stride of 2. (b) Max pooling with a pooling area of size 2x2 and stride of 2.
Global Pooling is a variant of pooling operation where the kernel size is equal to the size of the input feature map. Essentially, it computes a single value for each feature channel by summarizing the entire feature map. The two most common types of global pooling are global average pooling (GAP) and global max pooling (GMP), as illustrated in Fig. 2, which compute the average value and maximum value of each feature channel respectively. Global pooling is often used as a bridge between the backbone model and fully connected layers in classical image classification architectures. They help to reduce overfitting by minimizing the total number of parameters in the model. However, they can also lead to substantial loss of spatial information and critical details by transforming an entire feature map to a single value. This prevents the model from capturing local details effectively and thus limits the model performance.
Figure 2: Examples of two types of global pooling operations. (a) Global average pooling. (b) Global max pooling.
2.1.2 Quantum Basics
Here, we provide a brief introduction to some fundamental concepts of quantum computing that are essential for understanding this paper.
  • Qubit and Quantum State: The qubit is the basic unit of information in quantum computing. Unlike a classical bit which has a value of either 0 or 1, a qubit can exist in a combination of the two states. This is called superposition. The quantum state of a qubit is described by a vector in a two-dimensional Hilbert space, commonly represented as
    $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle,$ (1)
    where $|0\rangle$ and $|1\rangle$ are two computational basis states,  and $\alpha$ and $\beta$ are complex probability amplitude corresponding to each basis state, satisfying
    $|\alpha|^{2}+|\beta|^{2}=1.$
  • Quantum gate and Quantum circuit: Quantum gates are unitary operators or equivalently unitary matrices which are applied to states of qubits and perform unitary transformation. Common single-qubit quantum gates include the Hadamard gate, Pauli gates and their corresponding rotation gates. In particular, single-qubit rotation gates are parametric gates that depend on one parameter and allow for a more flexible range of operations. These rotation gates, denoted by $R_{i}(\theta)$, perform a rotation around the $i$ axis on the Bloch sphere [bloch1946nuclear] by an angle $\theta$, where $i\in\{x,y,z\}$ and $\theta\in[0,2\pi)$. Other examples of common gates include two-qubit gates such as the CNOT gate which can induce quantum entanglement between qubits. A summary of these quantum gates can be found in Table LABEL:gates.